OpenStudy (anonymous):

solve and simplify 0.4x+5<=1.2x-8

OpenStudy (angela210793):

0.4x-1.2x<=-13 -0.8x<=-13 x>=13/0.8

OpenStudy (amistre64):

you appear to be having the same kind of question over and over again ....

OpenStudy (anonymous):

4x+50<=12x-80 -8x<=130 x>=130/(-8)

OpenStudy (amistre64):

this is why the process is more important than any given answer ...

OpenStudy (anonymous):

the answer will be : x>=16.25

OpenStudy (anonymous):

\[x \ge \frac{130}{8}\]